OpenStudy (anonymous):

solve the equation: 70z^2-5z-15

hero (hero):

First step factor out 5 from each term

OpenStudy (anonymous):

5(14z^2 -z -3)

OpenStudy (anonymous):

can factor but without an= something, can not solve

OpenStudy (anonymous):

quadratic formula. (-b+/- sqrt(b^2-4ac))/2a x= .5 or x=-.4285

OpenStudy (anonymous):

sorry it equals 0

OpenStudy (anonymous):

To clarify the formula... you have it set up as... Az^2+Bz+C=0 where A B and C are all constants 70z^2+(-5)z+(-15)=0

OpenStudy (anonymous):

then you take out the five and set up parenthesis but then what

OpenStudy (anonymous):

so it's set up...(-(-5)+/- sqrt((-5^2)-4(70)(-15))/(2*70) the +/- means you do the equation once with a positive sign and a second time with a negative sign there. Does that make sense?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

You're welcome :D

hero (hero):

70z^2 - 5z - 15 5(14z^2-z - 3) 5(14z^2-7z + 6z -3) 5(7z(2z-1)+3(2z-1)) 5((2z-1)(7z+3)) z = 1/2, -3/7

hero (hero):

I hate when that happens.

OpenStudy (anonymous):

In my opinion Quadratic is easier :P

hero (hero):

No, you solved it before I did. I started solving it after you solved it already.

OpenStudy (anonymous):

Yes, i did. Look wayy up there.

hero (hero):

Yeah, but mine looks prettier

OpenStudy (anonymous):

Okay. :)

hero (hero):

And I solved it by factoring, showing that I'm not dependent on the quadratic formula, which to me, is always a last resort ( a resort I never use)

hero (hero):

I'm the master of solving algebra without quad formulas, FOIL, or traditional substitution or elimination methods

hero (hero):

sider 1993, any thoughts or comments?

OpenStudy (anonymous):

well if I had not taken calculus in high school before this college course of algebra I would have gotten lost with the quadratic equation method because of right now we are using what hero used.